OpenStudy (anonymous):

Simplify. x^8 ÷ x^4 · x^2 x^2?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

x^6

OpenStudy (anonymous):

a^m / a^n = a^(m-n)

OpenStudy (anonymous):

x^8 / x^4 = ???

OpenStudy (anonymous):

Then apply: \[\huge x^a x^b = x^{a+b}\]

OpenStudy (anonymous):

x^6?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

Welcome dear...